# Stratified Hazard Sampling: Minimal-Variance Event Scheduling for CTMC/DTMC Discrete Diffusion and Flow Models

### v1

@@ Abstract
CTMC/DTMC-based discrete generative models, including uniform-noise discrete diffusion (e.g., D3PM/CTDD) and discrete flow matching, enable non-autoregressive sequence generation by repeatedly replacing tokens through a time-inhomogeneous Markov process. Inference is typically implemented with step-based simulation: each token decides to jump via independent Bernoulli (or categorical) draws at every discretization step. Under uniform-noise initialization, where self-correction requires multiple edits per position, these independent decisions induce substantial variance in both the number and timing of edits, leading to characteristic failure modes such as under-editing (residual noise) or over-editing (cascading unnecessary substitutions), decreasing reproducibility.
We propose Stratified Hazard Sampling (SHS), a drop-in and hyperparameter-free inference principle for any sampler that admits a stay-vs.-replace decomposition. SHS models per-token edits as events driven by cumulative hazard (CTMC) or cumulative jump mass (DTMC) and places events by stratifying this cumulative quantity: with a single random phase per position, a token jumps whenever its accumulated hazard crosses unit-spaced thresholds. This preserves the expected number of jumps while achieving the minimum possible variance among unbiased integer estimators (bounded by $1/4$), without altering per-jump destination sampling and thus retaining multimodality. We also introduce a phase-allocation variant for blacklist-style lexical constraints that prioritizes early edits at high-risk positions to mitigate late-masking artifacts.
Machine Learning, ICML
@@ 1 Introduction
Non-autoregressive generation for high-dimensional discrete data (e.g., text, code) is a promising approach that can significantly reduce inference latency by updating all tokens in parallel. Recently, a growing family of methods has recast discrete generation as a time-inhomogeneous Markov process that transports an easy prior $p_{0}$ to the data distribution $p_{1}$. This includes discrete diffusion models defined on a discrete time grid (DTMC; e.g., D3PM), continuous-time Markov jump process formulations (CTMC; e.g., CTDD), and discrete flow matching variants. Despite differences in training objectives, their inference procedures share a common primitive: at each time (or step), each position either stays put or is replaced according to a learned categorical kernel, with a time-dependent mass/rate controlling how often replacements occur.
The core focus of this paper is on how the choice of initialization (prior) fundamentally alters the complexity and variance structure of the sampling trajectories. Consider mask-start initialization, also known as absorbing-state initialization (austin2021d3pm; campbell2022ctdd), which begins with all tokens as [MASK] and permits only unmasking: each position transitions monotonically from [MASK] to a token exactly once, resulting in highly simplistic sampling paths. This offers advantages in training and inference stability, as well as straightforward implementation. Yet, this monotonic unmasking-only structure inherently disallows natural re-jumps to “refine” already selected tokens. To enable revisions, one must insert a re-masking trick where we revert positions to [MASK] and then unmask again, which introduces inefficient extra transitions, model evaluations, and conflicts with the original monotonic CTMC design. In essence, this approach resembles an autoregressive (AR) process with parallel or randomized update orders, where each position is finalized only once.
In contrast, uniform-noise start (austin2021d3pm; lou2024sedd; schiff2025simple) (initializing each position with a random token) allows the timing and frequency of jumps at each position to be contextually determined based on the current sequence, naturally enabling multiple jumps. This provides a self-correction capability, where misplaced tokens can be iteratively refined to better fit the evolving context. However, this flexibility comes at the cost of increased trajectory complexity, amplifying uncertainties in jump counts and timings. In standard time-step approximations, this manifests as frequent failure modes at the tails. There are two common types of failure modes that occur. Under-editing (too few jumps) occurs when insufficient substitutions occur, leaving residual noise or local inconsistencies, and preventing full integration of contextual information before the process effectively halts. Conversely, over-editing (too many jumps) occurs when excessive substitutions cascade, disrupting even coherent segments and leading to repeated unnecessary edits, resulting in repetitions or distortions rather than convergence.
@@
The goal of this paper is to preserve the expressive power of multiple jumps (i.e., self-correction) inherent to uniform-noise-start CTMC discrete language models (LMs), while structurally eliminating unnecessary sampler variance—particularly the high variance in jump counts and event timings induced by the sampler. To achieve this, we propose Stratified Hazard Sampling (SHS), which models jumps as a non-homogeneous Poisson process (NHPP) and uses the cumulative hazard $S(t)=\int_{0}^{t}\lambda(s)\,ds$ to stratify event placement in this space. SHS triggers jumps when the cumulative hazard $S(t)$ crosses integer boundaries offset by a random $\theta\sim\mathrm{Uniform}(0,1)$. Unlike prior works (austin2021d3pm; campbell2022ctdd; lou2024sedd; schiff2025simple; gat2024dfm), which used probabilistic coin flips at fixed time steps to trigger the jumps, SHS preserves the expected jump count while bounding the jump count variance to a theoretical minimum (at most $1/4$;see Appendix B). Consequently, SHS mitigates the long-tail waiting times and “sampling luck” that are especially pronounced in uniform-noise starts, simultaneously suppressing under- and over-editing, and enabling more stable, reproducible sampling even with fewer neural function evaluations (NFEs).
In addition to variance reduction, we propose a simple yet effective lexically constrained safety mechanism built on our SHS approach. We focus on blacklist-based lexical constraints, where users explicitly specify a set of prohibited tokens and the sampler enforces this constraint via token-level filtering; this approach is attractive because it is transparent, controllable, and can provide hard safety guarantees without retraining (hokamp-liu-2017-lexically; post-vilar-2018-fast; hu-etal-2019-improved; gehman-etal-2020-realtoxicityprompts). However, naively masking and renormalizing prohibited tokens at each local transition generally deviates from the desired conditional distribution ($p(x\mid c)$), often causing “late masking” artifacts when high-risk positions are corrected only after other tokens have already adapted to an unsafe context. Leveraging SHS’s hazard-space event scheduling method described in the previous paragraph, we bias jump timing so that positions with high violation risk are more likely to jump earlier, thereby resolving unsafe candidates sooner and reducing the mismatch to the conditional target distribution.
@@ 2.1 CTMC/DTMC-based Discrete Generative Models
Let $\mathcal{V}$ be a vocabulary, $N$ the sequence length, and denote a sequence as $x=(x_{1},\dots,x_{N})\in\mathcal{V}^{N}$. We consider a continuous-time generative process $(X_{t})_{t\in[0,1]}$ with terminal distribution $p_{1}(x)$ and an initial distribution $p_{0}$ (often easy to sample, e.g., uniform or masked tokens). Many non-autoregressive discrete generative models can be expressed as a time-inhomogeneous Markov process on $\mathcal{V}^{N}$. We focus on the common single-site replacement structure (austin2021d3pm; campbell2022ctdd; gat2024dfm) where, at each (continuous or discrete) time, each position either stays unchanged or is replaced by a new token.
@@
A time-inhomogeneous continuous-time Markov chain (CTMC) is specified by a generator $Q_{t}$. For each position $i$ we define an escape rate $\lambda_{i}(t,x)\geq 0$ and a destination distribution $q_{t,i}(\cdot\mid x)\in\Delta(\mathcal{V})$. Let $x^{(i\leftarrow v)}$ denote the sequence obtained from $x$ by replacing $x_{i}$ with $v$. We parameterize off-diagonal rates as
@@
and set $Q_{t}(x\mid x)=-\sum_{i}\lambda_{i}(t,x)$. This local CTMC view subsumes continuous-time discrete diffusion models (e.g., CTDD) and flow-based formulations (e.g., DFM) whenever their one-step update admits a mixture-of-(stay vs. replace) forms.
@@
Under the standard step-based sampler, the total number of jumps at position $i$ is $J_{i}=\sum_{k=0}^{n-1}B_{ik}$, a Poisson-binomial random variable with $\mathbb{E}[J_{i}]=\sum_{k}p_{ik}$ and $\mathrm{Var}(J_{i})=\sum_{k}p_{ik}(1-p_{ik})$. This Poisson-binomial variance, amplified under uniform-noise initialization where multiple self-correction edits are required, is the sampler-induced variance targeted by SHS. For more general background on CTMCs, DTMCs, and their relationship, the reader is referred to standard probability references such as ross2023ipm and grimmett2020prp.
2.2 Lexical Safety as Conditioning, and Why Naive Masking Is Biased
We formalize token-level censorship (blacklists) as a hard lexical constraint. Let $\mathcal{B}\subset\mathcal{V}$ be a set of forbidden tokens and define the safe set of sequences
$\mathcal{A}\;=\;\{x\in\mathcal{V}^{N}:\forall i,\;x_{i}\notin\mathcal{B}\}.$ (4)
The distribution we ideally want is the conditional terminal distribution
$p_{1}(x\mid\mathcal{A})\;=\;\frac{p_{1}(x)\,\mathbf{1}[x\in\mathcal{A}]}{\mathbb{P}_{p_{1}}(X_{1}\in\mathcal{A})}.$ (5)
Exact conditioning via Doob $h$-transform.
For a Markov process, conditioning on a terminal event generally changes the dynamics at all intermediate times. Let
$h_{t}(x)\;=\;\mathbb{P}(X_{1}\in\mathcal{A}\mid X_{t}=x)$ (6)
be the probability of eventual safety from state $x$ at time $t\in[0,1)$. The generator of the conditioned Markov dynamics can be expressed via a Doob $h$-transform (doob1957):
and diagonal entries chosen so rows sum to zero. An analogous $h$-transform exists for DTMC transition kernels (rogers2000) and similarly alters both the step-wise move probabilities and the destination distribution via lookahead survival. Crucially, (7) reweights both (i) which transitions are taken and (ii) how frequently they are taken, according to global lookahead information encoded by $h_{t}$.
Naive masking creates error w.r.t. $p_{1}(x\mid\mathcal{A})$.
A common censorship heuristic (gehman-etal-2020-realtoxicityprompts) modifies only the destination distribution by
$q^{\textsc{mask}}_{t,i}(v\mid x)\;\propto\;q_{t,i}(v\mid x)\,\mathbf{1}[v\notin\mathcal{B}],$ (8)
and keeps the same escape rates $\lambda_{i}(t,x)$. This “naive masking” approach defines a different generator $Q_{t}^{\textsc{mask}}(x^{(i\leftarrow v)}\mid x)=\lambda_{i}(t,x)\,q^{\textsc{mask}}_{t,i}(v\mid x)$, which in general does not match the Doob-transformed generator (7).
Naive masking (8) effectively approximates the true lookahead weight $h_{t}(\cdot)$ by a binary surrogate that only checks whether a single-step proposal is forbidden. However, the correct conditional dynamics depends on how a proposed update changes the future probability of reaching $\mathcal{A}$ by time 1. Unless $h_{t}(y)/h_{t}(x)\equiv 1$ for all allowed transitions (a degenerate case), we have $Q_{t}^{\textsc{mask}}\neq Q_{t}^{(\mathcal{A})}$, and the induced terminal distribution under repeated masked updates is biased relative to (5).
The mismatch is especially pronounced in settings where the initial distribution $p_{0}$ assigns substantial mass to unsafe tokens (e.g., uniform-noise initialization). Then $h_{t}(x)$ varies widely across intermediate states, and the correct conditional process would preferentially accelerate transitions that remove high-risk tokens early. Naive masking alone cannot reproduce this global reweighting, since it does not model $h_{t}$.
@@ 3.1 Discrete Diffusion and CTMC Generative Models
Diffusion-style generative modeling has been extended to discrete spaces in several ways. Multinomial diffusion and related categorical constructions were studied by hoogeboom2021argmax. D3PM (austin2021d3pm) generalizes discrete diffusion by allowing structured transition matrices, including absorbing-state and nearest-neighbor corruptions, and demonstrates strong results on text and images. A fully continuous-time perspective that explicitly formulates both forward corruption and reverse generation as CTMCs was developed by campbell2022ctdd, connecting discrete diffusion sampling to Markov jump process simulation techniques. More recently, Score Entropy Discrete Diffusion (SEDD) (lou2024sedd) learns the reverse jump rates by estimating probability ratios (concrete scores) with a scalable score-entropy objective. In parallel, schiff2025simple revisited uniform-noise diffusion language models (UDLM), deriving discrete classifier-free/classifier-based guidance and improved variational training bounds, which are particularly relevant for controllable generation settings where multiple edits per position are common.
Beyond diffusion, flow-based viewpoints for discrete data have recently emerged. Discrete Flow Matching (DFM) (gat2024dfm) proposes a discrete analogue of flow matching for high-dimensional categorical data and reports strong performance in language and code generation. Other contemporaneous work explores geometric and manifold-aware discrete flow formulations.
@@ 3.2 Iterative Masked-Token Refinement for Non-Autoregressive Generation
A parallel line of work studies iterative refinement with masked language models. Mask-Predict (ghazvininejad2019maskpredict) performs parallel decoding by repeatedly masking and regenerating low-confidence tokens. MaskGIT (chang2022maskgit) introduces a confidence-based masking schedule for fast non-autoregressive synthesis, popularizing keep-the-most-confident refinement. For language modeling, several diffusion-inspired or simplex/continuous relaxations have been proposed, including DiffusionBERT (he2023diffusionbert), SSD-LM (han2022ssdlm), and continuous-time/continuous-space categorical diffusion formulations (dieleman2022cdcd).
3.3 Constrained and Safe Generation via Lexical Constraints
Controlling generation with explicit lexical constraints has a long history in decoding-time algorithms. Grid Beam Search (hokamp-liu-2017-lexically) and Dynamic Beam Allocation (post-vilar-2018-fast) enforce inclusion constraints efficiently, with later work improving batching and throughput (hu-etal-2019-improved). Finite-state and terminology-constrained decoding frameworks also appear in sequence generation and NMT settings.
In safety contexts, a common baseline is hard filtering or banning of disallowed tokens/phrases, but empirical evidence suggests that simple banned-word strategies are often insufficient as a complete safety solution, motivating principled approaches that better align with the desired conditional distribution (gehman-etal-2020-realtoxicityprompts). Our work focuses on the distributional mismatch induced by naive censorship and how to reduce it within CTMC-based discrete generative modeling.

### v2

@@ Abstract
Uniform-noise discrete diffusion and flow models (e.g., D3PM, SEDD, UDLM, DFM) generate sequences non-autoregressively by iteratively refining randomly initialized vocabulary tokens through multiple context-dependent replacements. These models are typically formulated as time-inhomogeneous CTMC/DTMC processes and sampled using independent Bernoulli change decisions at each discretization step. This induces Poisson-binomial variance in per-position jump counts that grows with the number of required edits, leading to the characteristic under-editing (residual noise) and over-editing (cascading substitutions) failure modes that degrade sample quality, especially under tight discretization budgets. In contrast, absorbing-state (mask-start) models avoid this instability by allowing each position to jump at most once.
We propose Stratified Hazard Sampling (SHS), a training-free, drop-in, and hyperparameter-free inference principle for any sampler that admits a stay-vs.-replace decomposition. SHS models per-token edits as events driven by cumulative hazard (CTMC) or cumulative jump mass (DTMC) and places events by stratifying this cumulative quantity: with a single random phase per position, a token is updated whenever its accumulated hazard crosses unit-spaced thresholds. This preserves the expected number of jumps while achieving the minimum possible conditional variance among unbiased integer estimators (bounded by $1/4$ for any fixed cumulative mass), without altering per-jump destination sampling and thus retaining multimodality. Experiments on uniform-noise discrete diffusion language models show that SHS consistently improves sample quality. We further show that SHS improves robustness under token-level blacklist filtering, with benefits increasing as lexical constraints grow more severe. Code is available at the git hub.
@@ 1 Introduction
Non-autoregressive generation for high-dimensional discrete data (e.g., text, code) is a promising approach that can significantly reduce inference latency by updating all tokens in parallel. A growing family of methods recasts discrete generation as a time-inhomogeneous Markov process (CTMC or DTMC) that transports an easy prior $p_{0}$ to the data distribution $p_{1}$, including discrete diffusion models (D3PM, SEDD, UDLM) and discrete flow matching (DFM). Among initialization strategies, absorbing-state (mask-start) methods (Austin et al., 2021; Campbell et al., 2022)—where each position unmasks exactly once—yield simple, low-variance trajectories but reduce generation to any-order parallel unmasking without iterative refinement. In contrast, uniform-noise initialization (Austin et al., 2021; Lou et al., 2024; Schiff et al., 2025) allows each position to undergo multiple context-dependent jumps, providing genuine self-correction capability at the cost of increased trajectory complexity.
However, this multi-jump flexibility comes at a practical cost in the standard step-based sampler. At each discretization step, every position independently decides whether to jump via a Bernoulli (or categorical) draw. The total number of edits per position is therefore a Poisson-binomial random variable whose variance $\sum_{k}p_{ik}(1-p_{ik})$ can grow linearly with the cumulative jump mass—precisely the regime where uniform-noise models operate, since meaningful generation typically requires multiple self-correction edits per position. This sampler-induced variance produces two characteristic failure modes at the tails of the jump-count distribution. Under-editing (too few jumps) occurs when insufficient substitutions leave residual noise or local inconsistencies, preventing full integration of contextual information before the process terminates. Conversely, over-editing (too many jumps) occurs when excessive substitutions cascade, disrupting even coherent segments and producing repetitions or distortions. These failure modes are not inherent to the uniform-noise formulation itself but are artifacts of the independent per-step sampling mechanism, and they worsen as the discretization budget (NFE) decreases.
@@
The goal of this paper is to preserve the expressive multi-jump dynamics of uniform-noise-start discrete generative models while structurally eliminating this unnecessary sampler variance. To this end, we propose Stratified Hazard Sampling (SHS), which requires no retraining, no additional hyperparameters, and no architectural modification—only a change to the inference-time sampling rule. SHS leverages the cumulative hazard of a non-homogeneous Poisson process (NHPP) and uses the cumulative hazard $S(t)=\int_{0}^{t}\lambda(s)\,ds$ to stratify event placement in this space. SHS triggers jumps when the cumulative hazard $S(t)$ crosses integer boundaries offset by a random $\theta\sim\mathrm{Uniform}(0,1)$. Unlike prior works (Austin et al., 2021; Campbell et al., 2022; Lou et al., 2024; Schiff et al., 2025; Gat et al., 2024), which used probabilistic coin flips at fixed time steps to trigger the jumps, SHS preserves the expected jump count while bounding the jump count variance to a theoretical minimum (at most $1/4$; see Appendix C). Consequently, SHS mitigates the long-tail waiting times and “sampling luck” that are especially pronounced in uniform-noise starts, simultaneously suppressing under- and over-editing, and enabling more stable, reproducible sampling even with fewer neural function evaluations (NFEs).
Token-level blacklists provide a simple and transparent form of lexical control: at each replacement, forbidden tokens are filtered out and the remaining distribution is renormalized. While widely used, this filtering makes sampling trajectories more brittle: a single position that remains unedited can dominate lexical metrics, and this brittleness worsens as the blacklist becomes more restrictive. In this work, we keep the filtering rule fixed and isolate the effect of event scheduling, showing that SHS yields substantially improved robustness.
@@ 2.1 CTMC/DTMC-based Discrete Generative Models
Let $\mathcal{V}$ be a vocabulary, $N$ the sequence length, and denote a sequence as $x=(x_{1},\dots,x_{N})\in\mathcal{V}^{N}$. We consider a continuous-time generative process $(X_{t})_{t\in[0,1]}$ with terminal distribution $p_{1}(x)$ and an initial distribution $p_{0}$ (often easy to sample, e.g., uniform or masked tokens).
Many non-autoregressive discrete generative models can be expressed as a time-inhomogeneous Markov process on $\mathcal{V}^{N}$. We focus on the common single-site replacement structure (Austin et al., 2021; Campbell et al., 2022; Gat et al., 2024) where, at each (continuous or discrete) time, each position either stays unchanged or is replaced by a new token.
@@
A time-inhomogeneous continuous-time Markov chain (CTMC) is specified by a generator $Q_{t}$. For each position $i$ we define an escape rate $\lambda_{i}(t,x)\geq 0$ and a conditional destination distribution $q_{t,i}(\cdot\mid x)\in\Delta(\mathcal{V})$ satisfying $q_{t,i}(x_{i}\mid x)=0$ (i.e., $q_{t,i}(\cdot\mid x)\in\Delta(\mathcal{V}\setminus\{x_{i}\})$), where $\Delta(\mathcal{V})$ denotes the probability simplex over $\mathcal{V}$. Let $x^{(i\leftarrow v)}$ denote the sequence obtained from $x$ by replacing $x_{i}$ with $v$. We parameterize off-diagonal rates as
@@
and set $Q_{t}(x\mid x)=-\sum_{i}\lambda_{i}(t,x)$. This local CTMC view subsumes continuous-time discrete diffusion models (e.g., CTDD) and flow-based formulations (e.g., DFM) whenever their one-step update admits a mixture-of-(stay vs. replace) form.
@@
Under the standard step-based sampler, the total number of jumps at position $i$ is $J_{i}=\sum_{k=0}^{n-1}B_{ik}$, a Poisson-binomial random variable with $\mathbb{E}[J_{i}]=\sum_{k}p_{ik}$ and $\mathrm{Var}(J_{i})=\sum_{k}p_{ik}(1-p_{ik})$. This Poisson-binomial variance, amplified under uniform-noise initialization where multiple self-correction edits are required, is the sampler-induced variance targeted by SHS. For more general background on CTMCs, DTMCs, and their relationship, the reader is referred to standard probability references such as Ross (2023) and Grimmett and Stirzaker (2020).
Absorbing-state (mask-start) vs. uniform-noise.
The sampler-induced jump-count variance above primarily matters in multi-jump (self-correction) regimes. In the standard absorbing-state (mask-start) setting, the reverse dynamics is unmask-only: each position transitions monotonically from [MASK] to a vocabulary token at most once, so the per-position jump count satisfies $J_{i}\in\{0,1\}$. See Appendix D.4 for a formal statement.
2.2 Lexical constraints via blacklist filtering
We use token-level blacklists as a simple lexical constraint and as a robustness stress test for step-based inference under uniform-noise initialization. Fix a blacklist ratio $\rho\in[0,1)$ and sample a forbidden set $\mathcal{B}_{\rho}\subset\mathcal{V}$ uniformly at random with $|\mathcal{B}_{\rho}|=\lfloor\rho|\mathcal{V}|\rfloor$. Let the allowed vocabulary be $\mathcal{V}_{\rho}:=\mathcal{V}\setminus\mathcal{B}_{\rho}$ and the allowed set of sequences $\mathcal{A}_{\rho}:=\{x\in\mathcal{V}^{N}:\forall i,\ x_{i}\in\mathcal{V}_{\rho}\}$.
Safe-vocabulary initialization.
To avoid conflating lexical filtering with explicit “token removal” dynamics, we initialize
$X_{0}\sim\mathrm{Unif}(\mathcal{V}_{\rho})^{N},$ (4)
i.e., each position starts from a uniformly random token in the allowed vocabulary.
Mass-preserving destination filtering (renormalization).
At inference time, we keep the model-predicted change mass (DTMC) $p_{ik}(x)$ in equation 2 (or the escape rate (CTMC) $\lambda_{i}(t,x)$ in equation 1) unchanged, and enforce the blacklist only by filtering and renormalizing the destination distribution:
$q^{(\rho)}_{t,i}(v\mid x)\;=\;\frac{q_{t,i}(v\mid x)\,\mathbf{1}[v\in\mathcal{V}_{\rho}]}{\sum_{u\in\mathcal{V}_{\rho}}q_{t,i}(u\mid x)}.$ (5)
This ensures that lexical filtering does not trivially reduce the expected number of effective edits by shrinking the total jump mass; it only restricts which tokens can be proposed when a jump occurs. In all blacklist experiments (Section 5.2), we apply the same filtering rule equation 5 to both the Standard sampler and SHS, and isolate the effect of event scheduling.
Conditioning lens (background).
One may view $\mathcal{A}_{\rho}$ as a terminal constraint and ask for the conditional distribution $p_{1}(\cdot\mid\mathcal{A}_{\rho})$. In general, the exact conditioned CTMC/DTMC dynamics corresponds to a Doob $h$-transform and can reweight not only destinations but also effective jump intensities, so masking-only destination filtering is generally biased as a conditional sampler. We use this only as a diagnostic lens and do not attempt to estimate $h_{t}$; see Appendix E for details.
@@ 3.1 Discrete Diffusion and CTMC Generative Models
Diffusion-style generative modeling has been extended to discrete spaces in several ways. Multinomial diffusion and related categorical constructions were studied by Hoogeboom et al. (2021). D3PM (Austin et al., 2021) generalizes discrete diffusion by allowing structured transition matrices, including absorbing-state and nearest-neighbor corruptions, and demonstrates strong results on text and images. A fully continuous-time perspective that explicitly formulates both forward corruption and reverse generation as CTMCs was developed by Campbell et al. (2022), connecting discrete diffusion sampling to Markov jump process simulation techniques. More recently, Score Entropy Discrete Diffusion (SEDD) (Lou et al., 2024) learns the reverse jump rates by estimating probability ratios (concrete scores) with a scalable score-entropy objective. In parallel, Schiff et al. (2025) revisited uniform-noise diffusion language models (UDLM), deriving discrete classifier-free/classifier-based guidance and improved variational training bounds, which are particularly relevant for controllable generation settings where multiple edits per position are common. Relatedly, von Rütte et al. (2025a) propose Generalized Interpolating Discrete Diffusion (GIDD), which generalizes masked diffusion by interpolating between data and a mixing distribution and explores hybrid masking/uniform-noise noising schemes. A systematic study of scaling behavior across noise types by von Rütte et al. (2025b) finds that uniform diffusion requires more parameters but less data for compute-efficient training, and scales a uniform diffusion model to 3B and 10B parameters.
Beyond diffusion, flow-based viewpoints for discrete data have recently emerged. Discrete Flow Matching (DFM) (Gat et al., 2024) proposes a discrete analogue of flow matching for high-dimensional categorical data and reports strong performance in language and code generation. Other contemporaneous work explores geometric and manifold-aware discrete flow formulations.
@@ 3.2 Iterative Masked-Token Refinement for Non-Autoregressive Generation
A parallel line of work studies iterative refinement with masked language models. Mask-Predict (Ghazvininejad et al., 2019) performs parallel decoding by repeatedly masking and regenerating low-confidence tokens. MaskGIT (Chang et al., 2022) introduces a confidence-based masking schedule for fast non-autoregressive synthesis, popularizing keep-the-most-confident refinement. For language modeling, several diffusion-inspired or simplex/continuous relaxations have been proposed, including DiffusionBERT (He et al., 2023), SSD-LM (Han et al., 2022), and continuous-time/continuous-space categorical diffusion formulations (Dieleman et al., 2022).